Find the volume charge density associated with a field given by D= (xy²)i + (yx²)j + (z) C/m²


1
Expert's answer
2021-05-24T14:14:07-0400

The given electric field is associated with D ="(xy^{2}) \hat{i}+(yx^{2})\hat{j}+z\ c/m^{2}"


now , we have to find out volume charge density with the help of given field function , the relation between the Electric field and volume charge density is given by -


volume charge density is denoted by "\rho" = "{\nabla}.D"


"{\nabla}=\dfrac{\partial}{\partial x}\hat{i}+\dfrac{\partial}{\partial y}\hat{j}+\dfrac{\partial}{\partial z}\hat{k}"


"\rho" "=\dfrac{\partial(xy^{2})}{\partial x}\hat{i}+\dfrac{\partial(yx^{2})}{\partial y}\hat{j}+\dfrac{\partial(z)}{\partial z}\hat{k}"


"\rho=y^{2}\hat{i}+x^{2}\hat{j}+\hat{k}" "c\/m^{3}"